OpenStudy (anonymous):

Hello, I need help with factoring this equation. It's supposed to have three parentheses but I am only familiar with two (and still having trouble with that). Enter the dimensions of a rectangular box with a volume of 50b3 - 3 - 2b + 75b2.

OpenStudy (amistre64):

cubics have a formula similar to quadratics ...

OpenStudy (amistre64):

IF we try a more brute method ... just expanding a general cubic setup and comparing, it may be simple or not so simple to compare parts

OpenStudy (amistre64):

(x + a)(x + b)(x + c) (x^2 + 2ab x + ab)(x + c) x^3 + 2ab x^2 + ab x c x^2 + 2abc x + abc ------------------------------ x^3 +(2ab+c)x^2 + (ab+2abc)x + abc

OpenStudy (amistre64):

50b3 - 3 - 2b + 75b2 50b^3+ 75b^2 - 2b - 3 50 x^3+ 75 x^2 - 2 x - 3 x^3+ 75/50 x^2 - 2/50 x - 3/50 x^3 +(2ab+c)x^2 + (ab+2abc)x + abc

OpenStudy (amistre64):

or if we make them decimals ... *2/2 x^3+ 1.50 x^2 - .04 x - .06 x^3 +(2ab+c)x^2 + (ab+2abc)x + abc 2ab + c = 1.50 ab+2abc = .04 abc = -.06 3 equations in 3 unknowns to play with

OpenStudy (amistre64):

2ab + c = 1.50 ab+2abc = .04 abc = -.06 <-- ab+2(.06) = .04 ab = .04 - 2(.06) c = 1.50 - 2ab c = 1.50 - 2(.04 - 2(.06)) seens to wirk itself out

OpenStudy (amistre64):

*seems to work itself out lol

OpenStudy (amistre64):

and i dropped a negative but thats easily remedied ... should be -.06

OpenStudy (alekos):

Very good But how do we find a and b?

OpenStudy (amistre64):

since we know c, we replace it in the original setup and work it out for a and b, which im assuming plays nice since i havent tried it

OpenStudy (anonymous):

50b^3 + 75b^2 - 2b-3 Factor by grouping. = 25 b^2 (2b+3) - 1 (2b + 3) = (25 b^2 - 1) (2b+3) Now you can factor the first part because it's a difference of squares. = (5b - 1)(5b + 1)(2b+3)

OpenStudy (amistre64):

c = 1.18 ... 2ab + 1.18 = 1.50 ab+2ab(1.18) = .04 ab(1.18) = -.06 2ab = 1.50 - 1.18 ab = (1.50 - 1.18)/2 etc ...

OpenStudy (amistre64):

grouping seems to be nice .... i never was good at grouping for some reason lol

OpenStudy (amistre64):

my idea has a tendency to introduce errors since theres more work involved. jam333 did very good

OpenStudy (anonymous):

So if the volume = (5b-1)(5b+1)(2b+3) and you need to give dimensions for a possible rectangular box, let's try b=1. That leads to the sides being 4, 6 and 5. There will be many possible answers. When b=2, the box would have sides 9, 11 and 7.

OpenStudy (amistre64):

(x + a)(x + b)(x + c) (x^2 + 2ab x + ab)(x + c) ^^^^ one error i see in mine ... (a+b) x, its a perfect square that has a=b ugh

OpenStudy (anonymous):

Thank you everyone for your replies, I'm sorry I should've mentioned I was trying to factor by grouping :)

OpenStudy (amistre64):

yes, that would have been some nice information to have lol